25) Your firm is one of the largest bakery’s in the area. As part of your risk management process, you are considering using options to hedge the price risk on your biggest input – wheat. You have determined that a price of R52/per ton would allow for you to keep the same profit margin as last year. The following wheat options offer a strike price of R50/per ton expiring in 1 month:
 Call options on wheat are selling at a premium of R0.87 per ton.
 Put options on wheat are selling for R0.72 per ton.
(a) Given the information above, will you need a call or a put option?
(b) If each option is for 100 tons, and you require 1000 tons of wheat, demonstrate the outcome if, at expiry, the spot price of wheat is (i) R40 per ton and (ii) R60 per ton.
1
Expert's answer
2020-07-09T14:29:03-0400

a)Call options: They provide the contract holder with an option to buy but not an obligation the underlying asset at a predetermined price in the future called the strike price at a fixed date.The company must therefore opt for the call option, so that it can buy wheat at the strike price of the option in future.

b) "(40-50)\times1000=-10 000"

"-10 000-0.87\times1000-10 000=-10870"


"(60-50)\times1000=10 000"

"10 000-0.87\times1000=9130"


The call option is exercised if the spot price of the underlying asset by the time the contract expires is higher execution price, and not executed if it is equal to or lower than the execution price